### v2

@@ 4 Experiments
4.1 Training & Investigating Five-Digit Addition Model
Training the model. The 5-digit 1-layer addition model in Quirke & Barez (2024) achieved an accuracy of approximately 99%. Our preliminary experiments indicated that a 2-layer, 3-head model was the smallest configuration capable of achieving fully accurate addition (see App. C for alternative configurations tested).
@@ 4.1 Training & Investigating Five-Digit Addition Model
This configuration effectively doubled the computational power compared to the 1-layer model. Moreover, a 2-layer model introduces the capability to “compose” the attention heads in novel ways, facilitating the implementation of more complex algorithms (Elhage et al., 2021b).
By enriching the training dataset with more instances of rare edge cases and using 30 thousand training batches, the 5-digit 2-layer model attained a final training loss of approximately $2.3\times 10^{-8}$. This model demonstrated a 1M Q accuracy (details in App. B).
Investigating the model. To gain insights into the model’s predictive behavior, we conducted an investigation focusing on its activity during prediction calculations. Ablation experiments targeting the nodes revealed that the model primarily depends on nodes located in token positions D’2 to A1, as illustrated in Figs 4 and 5. Across these nine positions, the model engages 36 nodes, comprising 21 attention head nodes and 15 MLP layer nodes. The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 6 and 7), elucidating the specific computations performed at each node.
@@
For each answer digit $A_{n}$using test questions corresponding to the T8, T9 and T10 categories, we performed Principal Component Analysis (PCA) on the nodes yielding interpretable results. Specifically nine “node and answer-digit” combinations (see Figs 13 and 14) showed strong clustering of the questions aligned to the T8, T9 and T10 categories, suggesting these nodes are implementing the TriCase categorisation.
Figure 4: This map shows the % of enriched questions that fail when we ablate each node in the 5-digit 2-layer 3-head addition model. The model only uses nodes in token positions P8 to P16 (tokens D’2 to A1). No attention heads in layer 1 are used. Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not used by the model.
Figure 5: This map shows the input tokens each attention head attends to at each token position. At token position P11 (token =) the model predicts the first answer digit A5. All digit pairs (e.g. D2 D’2) are attended to by P11.
Figure 6: This map shows the lowest complexity quanta S0, S1, etc impacted when we ablate each node in the 5-digit 2-layer 3-head addition model. To answer S0 questions, only the S0 nodes are used. To answer S1 questions, S0 and S1 nodes are used, etc.
Figure 7: This map shows the answer digit(s) A0 .. A5 impacted when we ablate each useful node in the 5-digit 2-layer 3-head addition model. The “column per digit” pattern in the last 6 columns mirrors the 1-layer model’s pattern. Cells containing values like A2..5 (in column D’0) are used in multiple prediction steps to calculate multiple answer digits e.g. A2 to A5.
@@
The algorithm predicts the first answer digit, A5, at position P11 (tokne =). This digit, which can only be 0 or 1, is the most challenging to predict as it may rely on a cascade starting from the digit pair D0 D’0 (e.g., 55555+44445=100000). The algorithm is required to compute this cascade using the nodes located in positions P8 to P11 (input tokens D’2 to =). As illustrated in Figure 5, these nodes attend to all digit pairs from D4 D’4 to D0 D’0 by P11. Additionally, the PCA data, as shown in Figures 13 and 14, suggest that these nodes produce tri-state outputs.
After encountering several initial setbacks (see App. E and F), we discovered that the model utilizes a minimal set of “make carry” information, leading to the development of the TriCase quanta.
The model performs $D_{n}.C$ using bigrams (see App. D) to map two input tokens to one result token e.g. “6” + “7” = T10. In positions P8 to P11, the model does $D_{n}.C$ calculations on all digit pairs from D4 D’4 to D0 D’0.
The mathematical framework behind this is intricate. To validate it, we implemented the algorithm in Python and successfully executed one million additions with zero errors.
@@
For a specific instance of the 5-digit addition model, we mapped the algorithm to individual nodes. Using ablation interventions, we verified each node’s role in executing the expected algorithmic step (see App. G). Table 2 and Figure 1 illustrate how the algorithm was applied to this particular model instance, adhering to all known constraints and accurately calculating all six answer digits.
Table 2: Pseudo-code for accurately predicting all answer digits, with mappings to the model nodes that implement the pseudo-code, for a specific 5-digit model instance.
Value Value calculated as Nodes used
D2.C
TriCase(D2, D’2)
P8.L0.H1 and P8.L0.MLP
D1.C
TriCase(D1, D’1)
P9.L0.H1 and P9.L0.MLP
D1.C2
TriAdd(D1.C, TriCase(D0, D’0))
P10.L0.H1 and P10.L0.MLP
D3.C4
TriAdd(TriCase(D3, D’3), TriAdd(D2.C,D1.C2))
P11.L0.H1
D4.C
TriCase(D4, D’4)
P11.L0.H2
D4.C5
TriAdd(D4.C, D3.C4)
P11.L0.MLP
A5
(D4.C5 == 10)
P11.L1.MLP
D4.BA
(D4 + D’4) % 10
P12.L0.H0 + H2
D3.C4
TriAdd(TriCase(D3, D’3), TriAdd(D2.C,D1.C2))
P12.L0.H1
A4
(D4.BA + (D3.C4 / 10)) % 10
P12.L0.MLP and P12.L1.MLP
D3.BA
(D3 + D’3) % 10
P13.L0.H0 + H2
D2.C3
TriAdd(D2.C,D1.C2)
P13.L0.H1
A3
(D3.BA + (D2.C3 / 10)) % 10
P13.L0.MLP and P13.L1.MLP
D2.BA
(D2 + D’2) % 10
P14.L0.H0 + H2
D1.C2
(D1 + D’1) / 10 + P10.D1.C2 *
P14.L0.H1
A2
(D2.BA + (D1.C2 / 10)) % 10
P14.L0.MLP and P14.L1.MLP
D1.BA
(D1 + D’1) % 10
P15.L0.H0 + H2
D0.MC
(D0 + D’0) / 10
P15.L0.H1
A1
(D1.BA + D0.MC) % 10
P15.L0.MLP and P15.L1.MLP
A0
(D0 + D’0) % 10
P16.L0.H0 + H2 P16.L0.MLP and P16.L1.MLP
Throughout the training of different instances of the 5-digit models, small modifications to the training environment resulted in subtly varied implementations. All variations seen use the Figure 3 algorithm, but each has small implementation differences. For example, in positions P8 to P11, the model might alter which 5 nodes perform the $D_{n}.C$ calculations and in what order. This corresponds to small changes in the Tab. 2 documentation per variation.
In summary, we concluded that the specific instance of the 5-digit addition model documented above is reliably accurate, well-understood, well-functioning and so known-good.
4.2 Case Study: Extending to Six-Digit Addition
Our investigations showed that all instances of the 5-digit addition model we trained appeared to utilize the same algorithm. To determine if this algorithm extends to related models, we trained a 6-digit 2-layer 3-head addition model, employing a different seed and optimizer. Additionally, we modified the answer format to include a ’+’ token (e.g., 111111+222222=+0333333).
Employing similar quanta and training data enrichment techniques, we developed a 6-digit addition model. This model achieved a loss of approximately $9.1\times 10^{-9}$ and demonstrated 1M Q accuracy.
The 5-digit (see Figures 4,  5,  6, and  7) and 6-digit (see Figures 15,  16,  8, and  17) quanta maps show numerous similarities, suggesting the 6-digit model implements the logical extension of the 5-digit algorithm (see Figure 3).
Intervention ablation tests show the 6-digit model does 6 BA and 6 MC calculations in late tokens, and does 6 $D_{n}.C$ calculations in early tokens with PCA showing the expected TriCase output. The 5 $D_{n}$.$C_{m}$ calculations are still to be tested.
Give this model has 1M Q accuracy, if this testing is successfully completed, and the algorithm implementation detail is documented, we can state the 6-digit model is known-good.
Figure 8: This map shows the simpliest addition complexity quanta S0, S1, etc used in each useful node of the 6-digit 2-layer 3-head addition model. Note the similarity to 5-digit Figure 6
4.3 Case Study: Extending the model for subtraction
To explore the reuse of known-good models, we developed a larger (6-digit 3-layer 4-head) model capable of performing both addition and subtraction. We introduced specific subtraction quanta and enriched the training dataset with rare subtraction edge cases.
This new “mixed” model, initially untrained, was initialized with the weights from the (6-digit 2-layer 3-head) addition model. To facilitate its learning of subtraction while maintaining its pre-inserted addition skills, we trained the mixed model with a dataset comprising 80% subtraction and 20% addition problems. This approach resulted in a model achieving 1M Q accuracy for both addition and subtraction tasks. Attempts to “freeze” useful attention heads and/or MLP layers from the addition model by periodically copying their weights back into the mixed model every 100 training epochs, however, led to decreased accuracy.
Subtraction ($a-b$) presents a greater challenge than addition ($a+b$), primarily due to its two distinct cases: $a>b$ (positive answer) and $a<b$ (negative answer). The model must discern this difference to accurately determine the first answer token as either “+” or “-”.
In cases of “positive” subtraction, the process closely mirrors addition. This observation led us to define subtraction quanta (see Tables 3 and 4 (see App. J), paralleling the addition quanta.
Calculating the “-” case in subtraction is more complex than the “+” case. The model might employ distinct circuits for each scenario. Our hypothesis is that after mastering the simpler “+” case, the model will learn to apply the mathematical identity $a-b=-(b-a)$, thus enabling it to handle “-” case questions by re-utilizing the “+” case circuits and then inverting the result. To examine this, we introduced the NG quanta (see Table 4) and again applied mean-ablation techniques to analyze the process.
Table 3: We categorise digit-pair differences (i.e. $D_{n}$- $D^{\prime}_{n}$ abbreviated as $D_{n}$.Diff) into three groups. These subtraction groups parallel the BA, MC and US addition groups. Groups BS and DZ overlap.
Name Abbrev Definition E.g. Like
Base Subtract BS $D_{n}$.Diff $<=0$ 7 - 2 BA
Borrow 1 B1 $D_{n}$.Diff $<0$ 7 - 9 MC
Diff Zero DZ $D_{n}$.Diff = 0 7 - 7 US
Table 4: We categorise subtraction questions into non-overlapping “calculation complexity” quanta, ordered by increased computational difficulty and decreasing occurrence frequency. Five-digit subtraction questions quanta are M0 to M4. Ten-digit subtraction question quanta are M0 to M9, and M9’s frequency is $\sim$0.000001%. Subtraction questions where the answer is negative get their own category NG .
Name Contains Example Freq
M0 BS 33333-22222=+11111 $\sim$25%
M1 BS,B1 33333-22282=+11051 $\sim$15%
M2 BS,B1,DZ 33333-22382=+10951 $\sim$8%
M3 BS,B1,DZx2 33333-23382=+09951 $\sim$2%
M4 BS,B1,DZx3 33333-23338=+09995 $\sim$0.1%
NG N/A 10000-90000=-80000 $\sim$50%
We compared the nodes utilized by the inserted addition model (see Figure 6) to those used by the mixed model when performing addition (see Figure 9). There is a strong correlation between them. Notably, the mixed model predominantly uses the existing attention heads in layers 1 and 2, ignoring many available spare heads.
Further comparison was made between the nodes the mixed model employs for addition and subtraction (see Figures 9 and 10). The nodes used for addition (S*) and those for “subtraction with a positive answer” (M*) show a very strong correlation. This suggests that the model repurposes its addition nodes to facilitate subtraction in some manner. Notably, an additional 15 nodes are activated for subtraction problems that result in a negative answer (represented by the NG quanta).
While the mixed model has 1M Q accuracy for both addition and subtraction tasks, intervation ablation testing and PCA of the mixed model have not been completed, and the algorithm implementation has not been documented. We cannot yet assert that it is known-good.
Figure 10: This map shows the simpliest complexity quanta M0, M1, etc used in each useful node of the 6-digit 3-layer 4-head mixed model for subtraction questions with positive answers. The NG cells are only used in subtractions $a-b$ where $a<b$, likely implementing the $-(b-a)$ identity as described in  4.3.

### v7

@@ 4 Experiments
4.1 Training a Five-Digit Addition Model
@@
The Quirke and Barez (2024) 5-digit 1-layer addition model achieved an accuracy of $\sim$ 99%. Our experiments suggested that a 2-layer, 3-head model was the smallest configuration capable of achieving 99.9999% accuracy (see App. D for alternatives tested). This configuration effectively doubled the computational power compared to the 1-layer model (see App. B for mode model configuration details). Moreover, a 2-layer model introduces the capability to “compose" the attention heads in novel ways, facilitating the implementation of more complex algorithms (Elhage et al., 2021b).
We trained a 5-digit, 2-layer, 3-head model, with a 14 token vocabulary ( 0, .., 9, +, -, =, *, / ), batch size of 64, learning rate of 0.00008 and weight decay of 0.1. Training used an infinite dataset enriched with rare edge cases. Loss was defined as the mean across all answer tokens of their negative log likelihood loss. After 30 thousand training batches, the model’s final training loss was $\sim 2.3\times 10^{-8}$. Testing showed this model has six nines accuracy. (More details in App. C and Tab. 6).
4.2 Investigating Five-Digit Addition
@@
Ablation experiments targeting the nodes revealed that the model depends only on nodes located in nine token positions (Figs 6 and 7). Further ablation experiments show that for these nine token positions, model uses 36 nodes in predictions. The effects of node ablation on our complexity and answer-impact metrics were analyzed (see Figs 5 and Table  1), providing insight into the specific computations performed at each node. For each answer digit $A_{n}$, using test questions corresponding to the ST8, ST9 and ST10 categories, we performed PCA on the nodes yielding interpretable results. Specifically nine “node and answer-digit” combinations (see Figs 8 and  9) showed strong clustering of the questions aligned to the ST8, ST9 and ST10 categories.
(P6) (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17)
D’4 D’1 D’0 = + A5 A4 A3 A2 A1
L0H0 A4 A2 A1
L0H1 A5 A5..3 A5 A3 A0
L0H2 A5 A5..1
L0MLP A5..2 A4 A3 A2 A1 A0
L1H0
L1H2 A4 A3 A2
L1MLP A1 A0
Table 1: The nodes used in predictions across token positions. For all addition and mixed models studied, before the "=" token, each node often calculates data used to predict multiple answer digits. After the "=" token, all nodes in a given token position are used to predict a single answer digit. For this sample model, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5.
@@
The algorithm predicts the first answer digit, A5, at position P11. A5, which is always 0 or 1, is the most challenging to predict as it may rely on a long carry one cascade (e.g. 55555+44445=100000). An accurate algorithm must compute this cascade using the nodes located in positions P8 to P11. As illustrated in Figure 7, these nodes attend to all digit pairs from D4 D’4 to D0 D’0. Additionally, the PCA data, as shown in Figures 8 and 9, suggest that these nodes produce tri-state outputs. After our first two algorithm hypothesises failed testing (see App. H and I), we discovered that the model utilizes a minimal set of “carry one” information, leading to the development of the TriCase quanta. The model performs $A_{n}.ST$ using bigrams (see App. F) to map two input tokens to one result token e.g. “6” + “7” = ST10. In positions P8 to P11, the model does $A_{n}.ST$ calculations on all digit pairs from D4 D’4 to D0 D’0.
An MLP layer can be thought of as a “key-value pair" memory (Meng et al., 2022; Geva et al., 2021) that can hold many bigrams and trigrams. We posit our MLP implements the TriAdd function using bigrams and trigrams to calculate $A_{n}.SV$ values from $A_{n}.ST$ values.
@@
For a specific 5-digit addition model instance, we mapped the algorithm to individual nodes. Using ablation intervention to override each node, we verified each node’s role in executing the expected algorithmic step. Figure 3 diagrams the addition algorithm, while App. J details each nodes purpose for this model instance. The model adheres to all known constraints. Empirically, it achieves six nines accuracy. In summary, we concluded that this specific instance of the 5-digit addition model documented above has six nines accuracy, is well-understood and well-functioning and hence known-good.
The 1-layer model uses 21 nodes to achieve two nines (99%) accuracy. This model uses 36 nodes (an increase of 71%) to achieve six nines accuracy.
4.3 Training n-digit Addition models
To investigate whether this algorithm is used widely, we first trained seven 2-layer addition models with 5-, 6- and 10-digits, using different seeds, a different optimizer, and changing the answer format to include a sign token (e.g. 111111+222222=+0333333).
These seven models all have very low loss (e.g. 1.5e-8) and six nines accuracy. (Details in Tab.6.)
4.4 Investigating n-digit Addition models
We implemented a declarative way to describe each sub-task. For example, a sub-task can be described as primarily attending to question digits D2 and D’2, influencing answer digit A3, impacting S0 but not S1 complexity questions, with a specific PCA result, and a specific ablation test. Using the declaration, we searched for and identified nodes performing all sub-tasks in all the models.
An algorithm hypothesis (e.g the one described in section 4.2) is then described in terms of the sub-tasks that must exist in the model and their inter-relationships. For example, our addition algorithm description states (in part) that the model must implement $A_{n}.ST$ sub-tasks for each question digit before the “=” token. We implemented an n-digit algorithm hypothesis declaration framework and a way to test the declaration against the sub-tasks found in each model and map the results. (Table 2 shows an example map.)
(P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
D’1 D’0 = + A6 A5 A4 A3 A2 A1
L0H0 A2.ST A3.ST A1.ST A4.ST A4.SC A3.SC A2.SC A1.SC A0.SC
L0H1 A1.ST A0.ST A5.SA A4.SA A3.SA A2.SA A1.SA A0.SA
L0H2 A5.ST
Table 2: All addition models studied implement our addition algorithm. The algorithm SA, SC and $ST$ sub-tasks all exist for each digit and in appropriate token positions. For a sample model, this map shows the subtask locations. Interestingly, here each SA sub-task is shared across two attention heads.
We found that our seven 2-layer addition models all implement our addition algorithm. Given these seven models have six nines accuracy, and all implement the same valid algorithm, we can state these models are known-good.
4.5 Training n-digit Mixed models
To explore reuse, we initialized untrained models with a known-good addition-only model, then trained them to perform both subtraction and addition. We call these “mixed" models.
Specifically, we trained seven larger (6- or 10-digit, 2- or 3-layer, 3- or 4-head) models after initializing them with the weights from a known-good 2-layer 3-head addition model. The first 2 layers, first 3 heads of the mixed model were initialized with the addition model weights. We trained the mixed model with 80% subtraction and 20% addition batches. We enriched the (infinite) training dataset with rare addition and subtraction edge cases.
Some models achieved six nines accuracy and the others five nines. Attempts to “freeze" the inserted attention heads and/or MLP layers by periodically copying the addition weights back into the mixed model every 100 training steps resulted in lower accuracy. (Refer App.M and Tab.6)
Figure 4: Each mixed model was initialized with the weights from a known-good addition model, then was trained on 80% subtraction and 20% addition batches. A sample log loss graph (final loss 8.0e-9) is shown.
4.6 Investigating n-digit Mixed models
Unlike addition, subtraction question answers can be either positive or negative. Similar looking positive-answer (e.g. 10009-10000=+000009) and negative-answer (e.g. 10009-20000=-009991) subtraction questions can give answers that differ at several digit positions. We posited that the model treats three distinct question classes “addition", “positive-answer subtraction" and “negative-answer subtraction" differently. Ablation of nodes showed that some nodes only help predict one question class, some help predict two classes and some all three classes. It also showed that the inserted addition nodes are heavily used and the majority become polysemantic, performing both addition and subtraction calculations. (Details in Tab.3.)
Used Inserted
Question class No % No %
All questions 96 48
Addition 61 64% 42 88%
Positive-answer sub 70 73% 40 83%
Negative-answer sub 53 55% 29 60%
Table 3: Mixed models re-use most inserted addition-model nodes. Many inserted nodes become polysemantic during training. That is they perform addition (S), positive-answer subtraction (M) and negative-answer subtraction (N) sub-tasks simultaneously. For a sample mixed model that uses 96 nodes and had 48 nodes inserted, this table shows inserted node reuse.
To investigate the subtraction algorithm, and paralleling addition algorithm subtasks, we defined subtraction-specific sub-tasks Base Diff and Borrow One (see Tab. 4) and complexity measures (see App. G and  N).
Name +ve Sub -ve Sub Definition
Base Diff MD ND $D_{n}$- $D^{\prime}_{n}$10
Borrow One MB NB $D_{n}$- $D^{\prime}_{n}$$<0$
Table 4: We define 2 “positive-answer subtraction" and 2 “negative-answer subtraction" sub-tasks that parallel the addition sub-tasks Base Add SA and Carry One SC
We found that the inserted nodes that performed SA in the addition model, perform SA, MD and ND in the mixed model. (Refer Table 5.) These three sub-tasks are similar in that each performs a mapping from 100 input cases (10 $D_{n}$x 10 $D^{\prime}_{n}$) to 10 output cases (0..9). The mixed model “upgraded" these nodes to be polysemantic during training. Similarly, some SC addition nodes became polysemantic and now process SC, MB and/or NB .
(P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
D’3 D’2 D’1 D’0 = A7 A6 A5 A4 A3 A2 A1
L0H0 A4.MT A3.MT A4.ST A2.SC A1.SC A0.SC OPR
A4.GT A3.MT A4.MT A4.SC A3.SC A2.NB A1.MB A0.MB SGN
OPR A1.NB A0.NB
L0H1 A2.ST A1.ST A3.ST A0.ST A5.SA
A4.ST A2.MT A1.MT A3.GT A0.MT A5.MD
A2.GT A1.GT A0.GT A4.SA A3.SA A2.SA A1.SA A0.SA
L0H2 A5.ST A5.SA A4.MD A3.MD A2.MD A1.MD A0.MD
OPR A5.MD A4.ND A3.ND A2.ND A1.ND A0.ND
SGN A5.ND
L0H3 OPR OPR OPR OPR OPR OPR
SGN SGN SGN SGN SGN SGN
Table 5: For mixed models, in later tokens, polysemantic attention heads simultaneously generate outputs for the three question classes addition S, M and N. Other heads calculate the question class by attending to the question operation (OPR) token and the answer sign (SGN) token. The MLP layers then select the output appropriate for the class. In this sample map, from P16, the first 3 rows contain many polysemantic nodes, while the 4th row calculates the question class.
An accurate subtraction model must answer “cascading borrow one" questions like 100000-000001=+0099999. We define the “essence of borrow one" subtask $MT$ used by both positive-answer (M) and negative answer (N) subtraction questions. $MT$ is like addition’s $ST$ subtask:
$\begin{split}A_{n}.MT&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}MTN&\text{if }D_{n}<D^{\prime}_{n}\\ MT0&\text{if }D_{n}=D^{\prime}_{n}\\ MT1&\text{if }D_{n}>D^{\prime}_{n}\\ \end{cases}\end{split}$ (5)
For M questions, MTN is a definite borrow one, MT0 is a possible borrow one (depending on the results of other digit calculations) and MT1 is definitely not a borrow one. For N questions, the interpretation is the opposite.
Paralleling addition’s “cascading carry one" $A_{n}.SV$ calculation, we define “cascading borrow one" calculation sub-tasks $A_{n}.MV$ for positive-answer subtraction and $A_{n}.NV$ for negative-answer subtraction. We posit the MLP implements TriAdd-like functions using bigrams and trigrams to calculate $A_{n}.MV$ and $A_{n}.NV$ values from $A_{n}.MT$ values.
4.7 Mixed model question class detection
We posit that there is a specific circuit to detect whether a question is in the S, M or N class. If the question operator is “+" then the class is S, but if the question operator is “-" then the model must calculate if $D>=D^{\prime}$ to distinguish between the M and N classes. For accuracy, it needs this answer by the “=" token to predict the first answer token (the answer sign) as “+" or “-".
$D>=D^{\prime}$ can be derived from the $A_{n}.MT$ data. Alternatively, this calculation could be a distinct circuit using a new sub-task we define as:
$A_{n}.GT=\begin{cases}1&\text{if }D_{n}\geq D^{\prime}_{n},\\ 0&\text{otherwise}.\end{cases}$ (6)
For a 4 digit question, $D>=D^{\prime}$ can be calculated as A3.GT=1 or (A3.GT=0 and (A2.GT=1 or (A2.GT=0 and (A1.GT=1 or (A1.GT=0 and A0.GT=1))). Our test for GT is that ablation causes the answer to change sign. The $MT$ calculation is very similar.
We found $A_{n}.MT$ and $A_{n}.GT$ subtasks in all mixed models. Usually (but not always) both subtasks are calculated by the same node. (Table 2 shows an example.) As both the $MT$ and GT approaches are valid, models can learn valid algorithm sub-task implementations that differ per answer digit.
We further posit that a node that implements say A2.SA, A2.MD and A2.ND does not know whether it is dealing with a S, M or N question, so it outputs all three possible answers to the residual stream. Another node must calculate the S, M or N distinction likely by attending to the question operator (OPR) and the answer sign (SGN). We define attention sub-tasks OPR and SGN for these calculations. Table 2 shows that for each answer digit one attention head attends to both OPR and SGN. We believe these heads transfer sufficient data to the residual stream to allow the MLP layer(s) to calculate the question class, and so select the appropriate output from the polysemantic A2.SA/MD/ND node.
4.8 Mixed model summary
The 7 mixed experimental models all have five or six nines accuracy. They all contain the same sub-tasks - suggesting they all implement the same algorithm. As we don’t fully understand the algorithm yet, we can’t assert that these mixed models are known-good.